# Spin-orbit interaction in tubular prismatic nanowires

### v1

@@ III.2 Nanowire states without SOI
Figure 5: Energy dispersions for the corner ($l=1,\dots 6$) and side ($l=7,\dots 12$) states of hexagonal shells with varying side thickness ($d$), where $R_{\mathrm{c}}=40$ nm and $r=2$ nm. The energies are shifted with respect to the ground energy at $k_{z}=0$. The line description defined on the right side of the figure applies for all panels. The label $l$ refers to a pair of degenerate states. In panels (b) and (c) one of the overlapping curves in each pair (representing levels: 2, 4, 8, and 10) is shown as dashed, with the dashed curve color corresponding to the same energy level as the solid line in panel (a).
Figure 6: Energy dispersions for the corner ($l=1,\dots 6$) and side ($l=7,\dots 12$) states of hexagonal shells with varying side thickness ($d$) and for the intermixing lengths equal to the side thicknesses ($r=d$). Here, $R_{\mathrm{c}}=40$ nm. The energies are shifted with respect to the ground energy at $k_{z}=0$. The line description defined on the right side of the figure applies for all panels. The label $l$ refers to a pair of degenerate states.
Now, we consider nanowires with infinite length. In the absence of SOI, the diagonalization of the Hamiltonian $\mathcal{H}_{cc}$ in Eq. (II) leads to parabolic energy dispersions, which reflect the degeneracy of the transverse states:
@@
For the 8-nm-thick hexagonal cross section [Fig. 3(a)], the corner and side states lead to the formation of eight parabolic curves, as shown in Fig. 4(a). Each group of states, corresponding to a particular localization, consists of twelve states arranged into four parabolas: the lowest parabola is always 2-fold degenerate (states $l=1$ and $l=5$), the next two curves are 4-fold degenerate (states $l=2$, $3$, $6$ and $l=7$), and the fourth parabola is again 2-fold degenerate (states $l=4$ and $l=8$).
In the case of the 8-nm-wide triangular and Tri-Hex shells shown in Fig. 1, the transverse corner states, visible in Figs. 3(b) and 3(c), are nearly degenerate. Consequently, for infinite-length wires, the corresponding energy levels form 6-fold quasi-degenerate parabolas, as shown in the lower panels of Figs. 4(b) and 4(c).
In the case of the triangular shell, the six states localized in the facets [Fig. 3(b)] can be described as follows: The first four transverse states, situated above the $\Delta_{\mathrm{Tri}}$ gap, form a 4-fold degenerate level, which gives rise to the 4-fold degenerate parabolic state ($l=3$) in Fig. 4(b). The remaining two side-localized states are 2-fold degenerate, resulting in double parabolic states ($l=4$) in the 3D structure. In contrast, for the Tri-Hex shell, the states above the $\Delta_{\mathrm{Tri-Hex}}$ gap are also localized in the corners [Fig. 3(c)]. Although these states are formally 4- and 2-fold degenerate, the energy separation between these levels is so small [Fig. 3(c)] that they almost align and when considering infinite shells, lead to 6-fold quasi-degenerate parabolic states ($l=3$ and $l=4$) in Fig. 4(c).
@@ III.3 Nanowire states with SOI
The different band gaps and affinities of the core and shell materials create a potential difference at the core-shell interface. This potential difference induces an electric field, which, in turn, gives rise to SOI. The potential profile can be modeled during the growth process [43] and can take a sharp, step-like shape or gradually decrease, i.e. can extend into the shell up to a distance $r$ (represented in dark blue in Fig. 1). Additionally, SOI can be induced by an external electric field which changes the shape of the confining potential, and thus creates structural asymmetry. The potential slope within the shell governs the Rashba coefficients which depend also on the energy gaps at the $\Gamma$ point of the shell [Eqs. (21) and (22)]. The effect of SOI on the energy bands depends primary on the fraction of the cross-sectional area where the Rashba coefficients are nonzero and on the overlap of this region with the electron localization areas.
@@ III.3.1 Hexagonal shell
In Fig. 6, we compare the energy dispersions of the corner and side states for symmetric hexagonal shells with varying side thicknesses. The shells surround identical cores ($R_{\mathrm{c}}=40$ nm) and the intermixing length ($r=2$ nm) is the same for each wire. The impact of SOI on the energy states is weak [Fig. 6(a)], and it decreases with increasing energy, due to the sharpening of the wave functions of the corner states and the appearance of smaller corner maxima in the side states’ group wave functions [53]. As energy increases, the shift of probability distribution towards the corner areas reduces the overlap between the wave functions and the region where the Rashba coefficients are nonzero (dark blue in Fig. 1).
@@
In the presence of SOI, the energy dispersions remain even functions of $k_{z}$. SOI lifts only the 4-fold degeneracies at $k_{z}\neq 0$, resulting in all states being 2-fold degenerate except at $k_{z}=0$ To avoid overlapping lines and double labeling, we label these pairs of states with $l$. Although the $z$ spin component ($\sigma_{z}$) is no longer a good quantum number, time reversal symmetry holds, so all these states are doubly degenerate via Kramers’ theorem. Nevertheless, the additional orbital degeneracy is still present at $k_{z}=0$, where SOI has no effect, and the energies are simply the transverse modes shown in Fig. 3(a) (before shifting the ground state to zero). This result is in disagreement with the cylindrical model of Rashba SOI, obtained by wrapping a planar electron gas on a cylinder [58], which lifts the orbital degeneracy at $k_{z}=0$ when imposed on a polygonal shell [12]. Still, our present result is in agreement with the $\bm{k}\!\cdot\!\bm{p}$ calculations for the core component of a core-shell nanowire [48].
@@
Figure 7: Energy dispersions for the corner and side (gray lines) states of the 8-nm-wide hexagonal shell in the presence of an external electric field. Here, $R_{\mathrm{c}}=40$ nm, $r=8$ nm, $E=0.1$ mV/nm, and $\varphi=\pi/12$. The energies are shifted with respect to the ground energy at $k_{z}=0$.
@@
As the shell thickness $d$ increases while the intermixing length $r$ is kept constant, the wave functions spread and the overlap with the region where the SOI coefficients are nonzero reduces. At the same time, the distinction between corner and side states vanishes, the energy gap between these groups of states decreases, and wide shells lose the properties of a prismatic structure, approaching the behavior of a cylindrical tube [52, 53]. The result is a gradual drop of the SOI effect, as seen in Figs. 6(b) and 6(c). In particular, for an 18-nm-wide shell [Fig. 6(c)], the energy dispersions resemble the parabolic form obtained in the absence of SOI [Fig. 4(a)].
@@
The impact of SOI is significantly amplified when the core material penetrates deeper into the shell. To illustrate this statement, in Fig. 6 we repeat the examples shown in Fig. 6, but for the intermixing length equal to the shell width ($r=d$). In this case, the band offset potential gradient extends from the core-shell interface to the external shell boundary and the SOI coefficients are nonzero over the entire cross section, maximizing the area on which the spatial derivatives in Eq. (21) are nonzero, and thus the SOI contribution to the Hamiltonian in Eq. (II). In addition, the tilted potential effectively narrows the shell by shifting the wave functions to the outer regions and modifying the electron distributions. As a result, for all three shell widths studied the SOI effects are more pronounced than in Fig. 6. In particular, the impact on the corner-localized states is significantly larger, the shift in the ±$\pm k_{z}$ directions is considerable, such that the lowest corner states become non-monotonic functions for both $k_{z}<0$ and $k_{z}>0$, the energy splittings at finite $k_{z}$ are more significant, and the states $l=3$ and $l=4$ cross.
@@
At larger $k_{z}$, the three lowest corner states change similarly with the wave vector along the growth direction. In contrast, the slopes of the three higher corner states ($l=4$, $5$, and $6$) decrease significantly with energy, leading to a reduction in the energy differences between them as $|k_{z}|$ increases, and thus to crossings. With the parameters used for Fig. 6(a), the SOI effect on the side-localized states is smaller than on corner states, causing the energy dispersions to remain almost parabolic functions of $k_{z}$. This is further due to the smaller overlap between the side-localized wave functions and the regions where the SOI coefficients are nonzero.
@@
Figure 8: Energy dispersions for the corner ($n=1,\dots 6$) and side ($n=7,\dots 12$) states of triangular shells with varying side thickness ($d$). Here, $R_{\mathrm{c}}=40$ nm and $r=2$ nm. The insets in panels (b) and (c) show a zoomed-in view of the corner states around $k_{z}=0$. The energies are shifted with respect to the ground energy at $k_{z}=0$. The line description defined on the right side of the figure applies for all panels.
@@
Figure 9: Energy dispersions for the corner ($n=1,\dots 6$) and side ($n=7,\dots 12$) states of the 8 and 18-nm-wide triangular shells with different intermixing lengths ($r$). The inset in panel (b) shows a zoomed-in view of the corner states around $k_{z}=0$. Here, $R_{\mathrm{c}}=40$ nm. The energies are shifted with respect to the ground energy at $k_{z}=0$. Line description defined on the right side of the figure is valid for all panels.
@@
For all studied wires, the potential drops from a value of $V_{\mathrm{CBO}}=656.6$ meV [44]. However, as the shell width increases while keeping $r=d$, the derivatives of the potential in Eqs. (21) decrease, leading to a reduced SOI contribution to the Hamiltonian in Eq. (II). Despite that the impact of SOI on the energy states of the 10-nm-wide shell [Fig. 6(b)] remains much stronger than when the potential penetrates only up to $r=2$ nm [Fig. 6(b)]. Although the potential derivatives in Eqs. (21) are more than half smaller for the 18-nm-wide shell [Fig. 6(c)] compared to the 8-nm shell, the effect of SOI on the energy states remains significant.
@@
In Fig. 7, we show the energy dispersions for the 8-nm-wide shell with $r=d$ [as in Fig. 6(a)] in the presence of a perpendicular electric field rotated by $\varphi=\pi/12$ with respect to the $x$-axis (Fig. 1). The contribution to SOI caused by the electric field is much smaller than that due to the extended band offset potential, but the impact of the field on the transverse energies and electron localization [45], and thus on the energy dispersions is considerable. The applied electric field lifts the 4-fold degeneracies at $k_{z}=0$ and the 2-fold degeneracies at finite $k_{z}$. As a result, the energies become 2-fold degenerate at $k_{z}=0$, and nondegenerate elsewhere. The corner states exhibit a spectrum resembling six slightly different wires, with six pairs of parabolas symmetrically shifted in the ±$\pm k_{z}$ direction. However, in contrast to independent systems, the corner states do not cross but repel at the otherwise crossing points. This level repulsion depends on the overlap of the wave function.
@@
The side states are more robust to an external electric field than the corner states [45]. However, these states are still affected by the electric field and form complex distributions that can consist of multiple maxima, which is reflected in the energy dispersion (gray lines in Fig. 7). The electric field induces energy gaps by lifting the 4-fold degeneracies and increases the energy splittings, particularly for the corner states, but considerably reduces the energy gap between the corner and side states at $k_{z}=0$, thereby decreasing the robustness of the corner states’ subspace.
@@
Realistic wires are never perfectly symmetric; geometric distortions, such as variations in the thickness of the shell facets, break the wave function symmetries in a manner similar to the effect of a perpendicular electric field [45, 25]. As a result, the energy dispersion in the presence of an electric field closely resembles the spectrum of technologically accessible structures. Moreover, since an electrically induced asymmetry can cause a geometrically symmetric tube to behave like a non-symmetric system, applying the electric field to geometrically distorted shells could, to some extent, restore the properties of symmetric structures.
@@ III.3.2 Triangular shell
Figure 10: Energy dispersions for the corner ($n=1,\dots 6$) and side ($n=7,\dots 12$) states of the 8-nm-wide triangular shell in the presence of an external electric field. Here, $R_{\mathrm{c}}=40$ nm, $r=d=8$ nm, $E=0.1$ mV/nm, $\varphi=\pi/6$. The energies are shifted with respect to the ground energy at $k_{z}=0$.
[…]
@@
[…]
@@
The effect of SOI strongly decreases with increasing shell width, as shown in Fig. 9(b) and Fig. 9(c). For wider shells the corner states are arranged in two distinct levels. The ground state remains 2-fold degenerate, while the 4-fold degenerate state splits into three states at finite $k_{z}$. As in the case of the hexagonal shell, when the shell thickness increases, the corner as well as side states merge and behave in the same manner. However, although the impact of SOI on the energies of the 18-nm triangular shell [Fig. 9(c)] is weak, it is still stronger than in the case of the 18-nm-wide hexagonal wire [Fig. 6(c)].
@@
In Fig. 9 we show the results when the radial potential gradient extends over a wider range within the shell. For the 8-nm shell, the energy splitting of the corner states at $k_{z}=2\cdot 10^{7}$ $\mathrm{m}^{-1}$ increases by approximately a factor of three [Figs. 9(a) vs. 9(a)]. Moreover, the side states become more similar to the corner states, and at finite $k_{z}$, they are arranged into two well-separated groups consisting of three closely spaced states [Fig. 9(a)]. Similarly to hexagonal wires, the increased SOI effect is related to the enlarged region with potential gradient and effectively increased confinement. Even if the InP molecules penetrate only up to half the width of the 18-nm-wide shell [Fig. 9(b)], the effects of SOI increase significantly with respect to the case when $r=2$ nm, and the energy dispersions resemble those of thinner shells [Fig. 9(a)]. For $r=d$, the energy dispersions of the 8-nm-wide [Fig. 9(a)] and 18-nm-thick [Fig. 9(c)] shells are qualitatively similar, though quantitatively weaker in the latter case.
@@
If a triangular wire is exposed to a perpendicular electric field rotated by an angle $\varphi=\pi/6$ with respect to the $x$ axis, the six transverse corner states are rearranged into three 2-fold degenerate levels. Each level corresponds to a probability distribution with a sharp maximum localized in a specific corner area [45]. Since the corner-localized maxima of the 8-nm-wide shell are separated by large areas from which electrons are depleted [Fig. 3(b)], the eigenstates associated with the three lowest energies do not overlap in the presence of the electric field. Therefore, in the low-energy regime, the triangular shell behaves as three independent wires with different ground energies. This behavior is illustrated in Fig. 10, the six lowest states form three pairs of parabolas shifted in ±$\pm k_{z}$ direction which cross at $k_{z}=0$ as well as at finite values of $k_{z}$.
@@
The side-localized wave functions, which form elongated maxima [Fig. 3(b)], are much more robust to the electric field than the corner states [45], and thus they overlap. Consequently, in the presence of SOI, the side states cross only at $k_{z}=0$ and repel each other for $k_{z}\neq 0$ when they come close. One special feature of the triangular structures is the large energy gap that protects the corner states from interactions with other states even in the presence of an electric field. This gap also allows for the manipulation of the corner state subspace; for example, by rotating the electric field, one can tune the energy splittings between the corner states at $k_{z}=0$ [59].

### v2

@@ III.2 Nanowire states without SOI
Figure 5: Energy dispersions for the corner ($l=1,\dots,6$) and side ($l=7,\dots,12$) states of hexagonal shells with varying side thicknesses ($d$), where $R_{\mathrm{c}}=40$ nm and $r=2$ nm. The energies are shifted with respect to the ground state energy at $k_{z}=0$. The line description defined on the right-hand side of the figure applies for all panels. The label $l$ refers to a pair of degenerate states. In (b) and (c) one of the overlapping curves in each pair (representing levels: 2, 4, 8, and 10) is shown as dashed, with the dashed curve color corresponding to the same energy level as the solid line in (a).
Figure 6: Energy dispersions for the corner ($l=1,\dots,6$) and side ($l=7,\dots,12$) states of hexagonal shells with varying side thicknesses ($d$) and for the intermixing lengths equal to the side thicknesses ($r=d$). Here, $R_{\mathrm{c}}=40$ nm. The energies are shifted with respect to the ground state energy at $k_{z}=0$. The line description defined on the right-hand side of the figure applies for all panels. The label $l$ refers to a pair of degenerate states.
Now, we consider nanowires with infinite length. In the absence of SOI, the diagonalization of the Hamiltonian $\mathcal{H}_{cc}$ in Eq. (II) leads to parabolic energy dispersions, which reflect the degeneracy of the transverse states
@@
For the 8-nm-thick hexagonal cross section [Fig. 3(a)], the corner and side states lead to the formation of eight parabolic curves, as shown in Fig. 4(a). Each group of states, corresponding to a particular localization, consists of 12 states arranged into four parabolas: The lowest parabola is always twofold degenerate (states $l=1$ and $l=5$), the next two curves are fourfold degenerate (states $l=2$, 33, 66 and $l=7$), and the fourth parabola is again twofold degenerate (states $l=4$ and $l=8$).
In the case of the 8-nm-wide triangular and Tri-Hex shells shown in Fig. 1, the transverse corner states, visible in Figs. 3(b) and 3(c), are nearly degenerate. Consequently, for infinite-length wires, the corresponding energy levels form sixfold quasidegenerate parabolas, as shown in the lower panels of Figs. 4(b) and 4(c).
In the case of the triangular shell, the six states localized in the facets [Fig. 3(b)] can be described as follows: The first four transverse states, situated above the $\Delta_{\mathrm{Tri}}$ gap, form a fourfold degenerate level, which gives rise to the fourfold degenerate parabolic state ($l=3$) in Fig. 4(b). The remaining two side-localized states are twofold degenerate, resulting in double parabolic states ($l=4$) in the 3D structure. In contrast, for the Tri-Hex shell, the states above the $\Delta_{\mathrm{Tri\textrm{-}Hex}}$ gap are also localized in the corners [Fig. 3(c)]. Although these states are formally four- and twofold degenerate, the energy separation between these levels is so small [Fig. 3(c)] that they almost align and when considering infinite shells, lead to sixfold quasidegenerate parabolic states ($l=3$ and $l=4$) in Fig. 4(c).
@@ III.3 Nanowire states with SOI
The different band gaps and affinities of the core and shell materials create a potential difference at the core-shell interface. This potential difference induces an electric field, which, in turn, gives rise to SOI. The potential profile can be modeled during the growth process [43] and can take a sharp, steplike shape or gradually decrease, i.e., can extend into the shell up to a distance $r$ (represented in dark blue in Fig. 1). Additionally, SOI can be induced by an external electric field which changes the shape of the confining potential, and thus creates structural asymmetry. The potential slope within the shell governs the Rashba coefficients which depend also on the energy gaps at the $\Gamma$ point of the shell [Eqs. (21) and (22)]. The effect of SOI on the energy bands depends primary on the fraction of the cross-sectional area where the Rashba coefficients are nonzero and on the overlap of this region with the electron localization areas.
@@ III.3.1 Hexagonal shell
In Fig. 6, we compare the energy dispersions of the corner and side states for symmetric hexagonal shells with varying side thicknesses. The shells surround identical cores ($R_{\mathrm{c}}=40$ nm) and the intermixing length ($r=2$ nm) is the same for each wire. The impact of SOI on the energy states is weak [Fig. 6(a)], and it decreases with increasing energy, due to the shift of probability distribution towards the corner areas [53], which reduces the overlap between the wave functions and the region where the Rashba coefficients are nonzero (dark blue in Fig. 1).
@@
In the presence of SOI, the energy dispersions remain even functions of $k_{z}$. SOI lifts only the fourfold degeneracies at $k_{z}\neq 0$, resulting in all states being twofold degenerate except at $k_{z}=0$. To avoid overlapping lines and double labeling, in this section we use $l$ to denote pairs of degenerate states. Although the $z$ spin component ($\sigma_{z}$) is no longer a good quantum number, time reversal symmetry holds, so all these states are doubly degenerate via Kramers’ theorem. Nevertheless, the additional orbital degeneracy is still present at $k_{z}=0$, where SOI has no effect, and the energies are simply the transverse modes shown in Fig. 3(a) (before shifting the ground state to zero). This result is in disagreement with the cylindrical model of Rashba SOI, obtained by wrapping a planar electron gas on a cylinder [58], which lifts the orbital degeneracy at $k_{z}=0$ when imposed on a polygonal shell [12]. Still, our present result is in agreement with the $\bm{k}\!\cdot\!\bm{p}$ calculations for the core component of a core-shell nanowire [48].
@@
Figure 7: Energy dispersions for the corner and side (gray lines) states of the 8-nm-wide hexagonal shell in the presence of an external electric field. Here, $R_{\mathrm{c}}=40$ nm, $r=8$ nm, $E=0.1$ mV/nm, and $\varphi=\pi/12$. The energies are shifted with respect to the ground state energy at $k_{z}=0$.
@@
As the shell thickness $d$ increases while the intermixing length $r$ is kept constant, the wave functions spread and the overlap with the region where the SOI coefficients are nonzero reduces. At the same time, the distinction between corner and side states vanishes and the energy gap between these groups of states decreases. Wide shells lose the properties of a prismatic structure, approaching the behavior of cylindrical tubes [52, 53]. The result is a gradual drop of the SOI effect, as seen in Figs. 6(b) and 6(c). In particular, for an 18-nm-wide shell [Fig. 6(c)], the energy dispersions resemble the parabolic form obtained in the absence of SOI [Fig. 4(a)].
@@
The impact of SOI is significantly amplified when the core material penetrates deeper into the shell. To illustrate this statement, in Fig. 6 we repeat the examples shown in Fig. 6, but for the intermixing length equal to the shell width ($r=d$). In this case, the band-offset potential gradient extends from the core-shell interface to the external shell boundary and the SOI coefficients are nonzero over the entire cross section, maximizing the area on which the spatial derivatives in Eq. (21) are nonzero as well as their values, and thus the SOI contribution to the Hamiltonian in Eq. (II). In addition, the tilted potential effectively narrows the shell by shifting the wave functions to the outer regions and modifying the electron distributions. As a result, for all three shell widths studied the SOI effects are more pronounced than in Fig. 6. In particular, the impact on the corner-localized states is significantly larger, the shift in the $\pm k_{z}$ directions is considerable, such that the lowest corner states become non-monotonic functions for both $k_{z}<0$ and $k_{z}>0$, the energy splittings at finite $k_{z}$ are more significant, and the states $l=3$ and $l=4$ cross at much smaller $k_{z}$.
@@
At larger $k_{z}$, the three lowest corner states change similarly with the wave vector along the growth direction. In contrast, the slopes of the three higher corner states ($l=4$, 55, and $6$) decrease significantly with energy, leading to a reduction in the energy differences between them as $|k_{z}|$ increases, and thus to crossings. With the parameters used for Fig. 6(a), the SOI effect on the side-localized states is smaller than on corner states, causing the energy dispersions to remain almost parabolic functions of $k_{z}$. This is due to the elongated shape of the side-localized maxima and their shift to regions where the extended band-offset potential is lower.
@@
Figure 8: Energy dispersions for the corner ($n=1,\dots,6$) and side ($n=7,\dots,12$) states of triangular shells with varying side thicknesses ($d$). Here, $R_{\mathrm{c}}=40$ nm and $r=2$ nm. The insets in (b) and (c) show a close-up view of the corner states around $k_{z}=0$. The energies are shifted with respect to the ground state energy at $k_{z}=0$. The line description defined on the right-hand side of the figure applies for all panels.
@@
Figure 9: Energy dispersions for the corner ($n=1,\dots,6$) and side ($n=7,\dots,12$) states of the 8 and 18-nm-wide triangular shells with different intermixing lengths ($r$). The inset in (b) shows a close-up view of the corner states around $k_{z}=0$. Here, $R_{\mathrm{c}}=40$ nm. The energies are shifted with respect to the ground state energy at $k_{z}=0$. The line description defined on the right-hand side of the figure is valid for all panels.
@@
For all studied wires, the potential drops from a value of $V_{\mathrm{CBO}}=656.6$ meV [44]. However, as the shell width increases while keeping $r=d$, the derivatives of the potential in Eqs. (21) decrease, leading to a reduced SOI contribution to the Hamiltonian in Eq. (II). Even so, that the impact of SOI on the energy states of the 10-nm-wide shell [Fig. 6(b)] remains much stronger than when the potential penetrates only up to $r=2$ nm [Fig. 6(b)]. Although the potential derivatives in Eqs. (21) are more than half smaller for the 18-nm-wide shell [Fig. 6(c)] compared to the 8-nm shell, the effect of SOI on the energy states remains significant.
@@
In Fig. 7, we show the energy dispersions for the 8-nm-wide shell with $r=d$ [as in Fig. 6(a)] in the presence of a perpendicular electric field rotated by $\varphi=\pi/12$ with respect to the $x$ axis (Fig. 1). The contribution to SOI caused by the electric field is much smaller than that due to the extended band-offset potential, but the impact of the field on the transverse energies and electron localization [45], and thus on the energy dispersions is considerable. The applied electric field lifts the fourfold degeneracies at $k_{z}=0$ and the twofold degeneracies at finite $k_{z}$. As a result, the energies become twofold degenerate at $k_{z}=0$, and nondegenerate elsewhere. The corner states exhibit a spectrum resembling six slightly different wires, with six pairs of parabolas symmetrically shifted in the $\pm k_{z}$ direction. However, in contrast to independent systems, the corner states do not cross but repel at the otherwise crossing points. This level repulsion depends on the overlap of the wave function and is the indication of an interaction.
@@
The side states are more robust to an external electric field than the corner states [45]. However, these states are still affected by the electric field and form complex distributions that can consist of multiple maxima, which is reflected in the energy dispersion (gray lines in Fig. 7). The electric field induces energy gaps by lifting the fourfold degeneracies and increases the energy splittings, particularly for the corner states, but considerably reduces the energy gap between the corner and side states at $k_{z}=0$, thereby decreasing the robustness of the corner states’ subspace.
@@
Realistic wires are never perfectly symmetric; Geometric distortions, such as variations in the thickness of the shell facets, break the wave function symmetries in a manner similar to the effect of a perpendicular electric field [45, 25]. As a result, the energy dispersion in the presence of an electric field closely resembles the spectrum of technologically accessible structures. Moreover, since an electrically induced asymmetry can cause a geometrically symmetric tube to behave as a nonsymmetric system, applying the electric field to geometrically distorted shells could, to some extent, restore the properties of symmetric structures.
@@ III.3.2 Triangular shell
Figure 10: Energy dispersions for the corner ($n=1,\dots,6$) and side ($n=7,\dots,12$) states of the 8-nm-wide triangular shell in the presence of an external electric field. Here, $R_{\mathrm{c}}=40$ nm, $r=d=8$ nm, $E=0.1$ mV/nm, $\varphi=\pi/6$. The energies are shifted with respect to the ground state energy at $k_{z}=0$.
The properties of quantum wires that originate from their polygonal cross sections, such as corner localization and the energy gap between the corner and side states, become more pronounced as the number of corners decreases. Thus these effects are the most significant for triangular structures [45]. In Fig. 9(a), we show the energy dispersion for an 8-nm-wide triangular shell. Although the ground transverse state (twofold degenerate) and the first excited transverse state (fourfold degenerate) are energetically resolved, they are very close to each other [Fig. 3(b)], and thus the corner-localized states can be considered a sixfold quasidegenerate level. The corresponding probability distributions of the corner states consist of three well-separated, narrow peaks, so the SOI affects these states as it would affect three independent wires. The corner states [states $n=1$ to $n=6$ in Fig. 9(a)] are thus arranged in three superimposed pairs, with the states shifted by wave vectors $\pm q=0.44\times 10^{7}$ m-1, which corresponds to $\alpha=\frac{\hbar^{2}}{m^{*}}q\approx 15$ meV nm when the approximation of shifted parabolas is used, i.e.
@@
The effect of SOI on the side states [from $n=7$ to $n=12$ in Fig. 9(a)] looks qualitatively different. In particular, the lower state, fourfold degenerate at $k_{z}=0$, splits into three states at finite $k_{z}$. The nondegenerate states ($n=7$ and $n=10$) arise from parabolas shifted by $\pm q\approx 0.88\times 10^{7}$ m-1 ($\alpha\approx 30$ meV nm), while the remaining two states ($n=8$ and $n=9$) form twofold quasiparabolic state with an effective mass that is larger than that predicted by Eq. (20). Similarly, the states $n=11$ and $n=12$ merge into one quasi-parabolic state, but with an effective mass smaller than that given in Eq. (20). Unlike a hexagonal cross section, the triangular cross section lacks inversion symmetry, and as a result, the degeneracies at $k_{z}\neq 0$ are partially lifted if the wave functions overlap. Also, electrons with energies above the $\Delta_{\mathrm{Tri}}$ gap, i.e., above the corner states, can occupy nearly the entire cross-sectional area, and thus the SOI affects these states differently than the states locked at corners.
@@
As in the case of hexagonal shells, the effect of SOI strongly decreases with increasing shell width when $r$ is kept constant [Figs. 9(b) and 9(c)]. For wider shells, the corner states are arranged into two distinct levels at $k_{z}=0$, the higher one splits into three states at finite $k_{z}$, thus behaving as the fourfold degenerate side-localized state. Although the impact of SOI on the energies of the 18-nm triangular shell [Fig. 9(c)] is weak, it is still stronger than in the case of the 18-nm-wide hexagonal wire [Fig. 6(c)].
@@
In Fig. 9 we show the results when the radial potential gradient extends over a wider range within the shell. For the 8-nm shell, the energy splitting of the corner states at $k_{z}=2\times 10^{7}$ $\mathrm{m}^{-1}$ increases by approximately a factor of three [Figs. 9(a) vs. 9(a)]. Moreover, the side states become more similar to the corner states, and at finite $k_{z}$, they are arranged into two well-separated groups consisting of three closely spaced states [Fig. 9(a)]. Similarly to hexagonal wires, the increased SOI effect is related to the enlarged region with potential gradient and effectively increased confinement. Even if the P atoms penetrate only up to half the width of the 18-nm-wide shell [Fig. 9(b)], the effects of SOI increase significantly with respect to the case when $r=2$ nm [Fig. 9(c)], and the energy dispersions resemble those of thinner shells [Fig. 9(a)]. For $r=d$, the energy dispersions of the 8-nm-wide [Fig. 9(a)] and 18-nm-thick [Fig. 9(c)] shells are qualitatively similar, though quantitatively weaker in the latter case.
@@
If a triangular wire is exposed to a perpendicular electric field rotated by an angle $\varphi=\pi/6$ with respect to the $x$ axis, the six transverse corner states are rearranged into three twofold degenerate levels. Each level corresponds to a probability distribution with a sharp maximum localized in a specific corner area [45]. Since the corner-localized maxima of the 8-nm-wide shell are separated by large areas from which electrons are depleted [Fig. 3(b)], the eigenstates associated with the three lowest energies do not overlap in the presence of the electric field. Therefore, in the low-energy regime, the triangular shell behaves as three independent wires with different ground energies. This behavior is illustrated in Fig. 10, where the six lowest states form three pairs of parabolas shifted in the $\pm k_{z}$ direction which cross at $k_{z}=0$ as well as at finite values of $k_{z}$, which confirms the absence of interaction.
@@
The side-localized wave functions, which form elongated maxima [Fig. 3(b)], are much more robust to the electric field than the corner states [45], and thus they overlap. Consequently, in the presence of SOI, the side states cross only at $k_{z}=0$ and repel each other for $k_{z}\neq 0$ when they come close which indicates the presence of interaction.
One special feature of the triangular structures is the large energy gap that protects the corner states from interactions with other states even in the presence of an electric field. This gap also allows for the manipulation of the corner state subspace; for example, by rotating the electric field, one can tune the energy splittings between the corner states at $k_{z}=0$ [59].